# Nanchong, Sichuan 637002, People’s Republic of China

### v1

@@ Abstract
Keywords:
Black Holes, Classical Theories of Gravity
1 Introduction
The search for quantum gravity effects has attracted attention over the last decades. On the experimental side, we have witnessed remarkable advancements, such as gravitational wave detections produced by LIGO and Virgo Abbott et al. (2016a, b, 2017, 2020), as well as high-energy particle collisions at the LHC Aad et al. (2012); Chatrchyan et al. (2012); Aaij et al. (2023). On the theoretical side, the study of Lorentz symmetry breaking (LSB) is important for understanding quantum gravity processes in fundamental physics Kostelecky and Potting (1991); Colladay and Kostelecky (1998); Kostelecky and Samuel (1989a); Gambini and Pullin (1999); Carroll et al. (2001); Tian and Du (2021); Tian et al. (2022). By investigating the low-energy contributions from Lorentz symmetry breaking, especially its impact on spacetime, and analyzing high-precision experimental data, we can explore the possibility of Lorentz symmetry breaking in spacetimes compared to general relativity (GR). This analysis also allows us to assess whether these implications are consistent with a quantum gravity model, providing an accurate description of the observed phenomena within that framework.
@@ 1 Introduction
Another significant Lorentz violation model is the Kalb-Ramond (KR) gravity theory, which incorporates a nonminimally coupled tensor field $B_{ab}$, known as the KR field. This field is a tensorial field arising from the bosonic spectrum of string theory Kalb and Ramond (1974); Kao et al. (1996). It is an antisymmetric 2-tensor, denoted as $B_{[ab]}$, and can be conveniently decomposed as $B_{ab}=\tilde{E}_{[a}v_{b]}+\epsilon_{abcd}v^{c}\tilde{B}^{d}$, where $v^{a}$ is a timelike 4-vector. The pseudofields $\tilde{E}_{a}$ and $\tilde{B}_{a}$ are spacelike, satisfying $\tilde{E}_{a}v^{a}=\tilde{B}_{a}v^{a}=0$. Analogous to Maxwell electrodynamics, these fields can be interpreted as the pseudoelectric and pseudomagnetic fields, respectively. Therefore, the KR VEV yields two background vectors, in contrast to the single vector produced by the bumblebee VEV. In this framework, considerable research has been conducted on finding exact solutions in the Einstein-Kalb-Ramond field equations Chakraborty and SenGupta (2016); Maluf and Muniz (2022); Yang et al. (2023); Duan et al. (2023). Notably, the pioneering work by Yang et al. Yang et al. (2023) first correctly derived the Schwarzschild-like solution in this theory, and the Schwarzschild-(A)dS-like solution was also obtained by relaxing the vacuum conditions. Subsequently, the study of spacetime properties surrounding these black hole solutions has garnered significant attention in the past year, including research on black hole quasi-normal mode (QNM) frequencies Guo et al. (2023); Filho et al. (2024a), photon spheres and shadows Jha (2024); Junior et al. (2024a), gravitational lensing Junior et al. (2024b), geodesic motion Filho et al. (2024b), etc Du et al. (2024). However, these solutions were limited to the specific spherically symmetric scenario $(-g_{tt}=g_{rr}^{-1})$ and did not extend to the more general spherically symmetric black hole scenario $(-g_{tt}\neq g_{rr}^{-1})$. To achieve this goal, it is essential to obtain a comprehensive understanding of solutions for static neutral spherically symmetric black holes solutions in KR gravity theory.
@@
The remaining part of this paper is organized as follows. In Sec. 2, we provide a concise overview of the Kalb-Ramond gravity theory. We solve the equations of the theory to obtain black hole exact solutions for two types of static neutral spacetimes (the cases where $-g_{tt}=g_{rr}^{-1}$ and $-g_{tt}\neq g_{rr}^{-1}$), considering both with and without the cosmological constant. In Sec. 3, some basic thermodynamic properties of the obtained Kalb-Ramond black holes are analyzed. In Sec. 4, we explore the dynamical evolution of the scalar field for both types of spacetimes. Both types of solutions decay rapidly with an increase in the Lorentz-violating parameter, with the latter type being more sensitive. Finally, our conclusions and outlooks are given in Sec. 5.
2 KR black hole solutions
@@
$T^{\text{KR}}_{ab}=$ $\frac{1}{2}H_{acd}H_{b}{}^{cd}-\frac{1}{12}g_{ab}H^{cde}H_{cde}+2V^{\prime}(X)B_{ca}B^{c}{}_{b}-g_{ab}V(X)+\xi_{2}\bigg{[}\frac{1}{2}g_{ab}B^{ce}B^{d}{}_{e}R_{cd}$ (3)
$-B^{c}{}_{a}B^{d}{}_{b}R_{cd}-B^{cd}B_{bd}R_{ac}-B^{cd}B_{ad}R_{bc}+\frac{1}{2}\nabla_{c}\nabla_{a}(B^{cd}B_{bd})+\frac{1}{2}\nabla_{c}\nabla_{b}(B^{cd}B_{ad})$
$-\frac{1}{2}\nabla^{c}\nabla_{c}(B_{a}{}^{d}B_{bd})-\frac{1}{2}g_{ab}\nabla_{c}\nabla_{d}(B^{ce}B^{d}{}_{e})\bigg{]},$
which can be considered as the energy-momentum tensor of the KR field. The prime indicates the derivative with respect to the variable of the corresponding functions.
@@
it follows that the vacuum field exhibits a pseudoelectric configuration. Consequently, this configuration automatically causes the KR field strength to vanish, i.e., $H_{abc}=0$ or $H_{3}=d\beta_{2}=0$ Yang et al. (2023). Then, we can define the efficient gravitational field equation, satisfying $\mathcal{G}_{ab}=0$, as follows:
$\mathcal{G}_{ab}=R_{ab}-\Lambda g_{ab}-\xi_{2}\mathcal{B}_{ab}-(2\beta_{ac}\beta_{b}{}^{c}+b^{2}g_{ab})V^{\prime},$ (4)
@@
$\mathcal{B}_{ab}=$ $g_{ab}\beta^{ce}\beta^{d}{}_{e}R_{cd}-\beta^{c}{}_{a}\beta^{d}{}_{b}R_{cd}-\beta^{cd}\beta_{ad}R_{bc}-\beta^{cd}\beta_{bd}R_{ac}+\frac{1}{2}\nabla_{c}\nabla_{a}(\beta^{cd}\beta_{bd})$ (5)
@@
and $\mathcal{G}_{\varphi\varphi}=\mathcal{G}_{\theta\theta}\sin^{2}\theta$, where $\varUpsilon\equiv(A^{\prime}/A)^{2}-2A^{\prime\prime}/A+A^{\prime}B^{\prime}/(AB)$ and $\ell\equiv\xi_{2}b^{2}/2$.
2.1 Asymptotically flat spacetime
We first consider asymptotically flat spacetimes, which do not include a cosmological constant, and aim to construct two different types of static neutral spherically symmetric black hole solutions within KR gravity. In this case, we take the assumption that $V^{\prime}=0$, which corresponds to the case where the VEV is located at the local minimum of the potential. A prime exemplar of a potential satisfying these conditions is a smooth quadratic function, similar to that assumed by Casana et al. Casana et al. (2018), denoted as:
$V=V(X)=\frac{1}{2}\lambda X^{2},$ (11)
where $\lambda$ serves as a constant, and $X$ represents a generic potential argument. Consequently, the VEV, $\beta_{ab}$, emerges as a solution of $V=V^{\prime}=0$.
Then, using Eqs. (8) and (9), we can obtain a simple relation:
@@ 2.1 Asymptotically flat spacetime
$ds^{2}_{\text{A}}=-\left(1-\frac{2M}{r}\right)dt^{2}+\frac{1-\ell}{\left(1-\frac{2M}{r}\right)}dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\varphi^{2}.$ (17)
It is straightforward to verify that the above solution (17) satisfies all field equations (4) when $\Lambda=V^{\prime}=0$. The dimensionless parameter $\ell$ represents the impact of Lorentz violation on spacetime, which is caused by the VEV of the KR field. Note that because the Riemann tensor is given by
@@
$ds^{2}_{\text{B}}=-\left(\frac{1}{1-\ell}-\frac{2M}{r}\right)dt^{2}+\frac{1}{\frac{1}{1-\ell}-\frac{2M}{r}}dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\varphi^{2}.$ (21)
@@
2.2 Asymptotically (A)dS spacetime
To investigate the effect of Lorentz violation in asymptotically (A)dS spacetime, that is, when a nonzero cosmological constant exists, consider relaxing the vacuum conditions. Similar to the potential given in Ref. Maluf and Neves (2021), another simple choice of potential is a linear function given as:
$V=V(\lambda,X)=\lambda X.$ (22)
In this context, $\lambda$ is interpreted as a Lagrange-multiplier field Bluhm et al. (2008). Its equation of motion ensures the vacuum condition $X=0$, leading to $V=0$ for any field $\lambda$ on-shell. Notably, this form of potential manifests itself as a cosmological constant. Thus, the potential and its first-order derivative satisfy
$V(\beta^{ab}\beta_{ab}+b^{2})=\lambda(\beta^{ab}\beta_{ab}+b^{2})=0,$ (23)
$V^{\prime}(\beta^{ab}\beta_{ab}+b^{2})=\lambda,$ (24)
where $V^{\prime}(X)=dV(X)/dX$.
It is evident that the relation given by Eq. (13) still holds. By setting the metric function $B(r)=\mathcal{C}_{1}A(r)^{-1}$, the equation induced by the relation $\mathcal{G}_{\theta\theta}=0$ is rewritten as
@@ 2.2 Asymptotically (A)dS spacetime
$ds^{2}=-\left[1-\frac{2M}{r}-\frac{(1-\ell)\Lambda_{e}}{3}r^{2}\right]dt^{2}+\frac{(1-\ell)dr^{2}}{1-\frac{2M}{r}-\frac{(1-\ell)\Lambda_{e}}{3}r^{2}}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\varphi^{2}.$ (29)
@@
$ds^{2}=-\left[\frac{1}{1-\ell}-\frac{2M}{r}-\frac{\Lambda_{e}}{3}r^{2}\right]dt^{2}+\frac{1}{\frac{1}{1-\ell}-\frac{2M}{r}-\frac{\Lambda_{e}}{3}r^{2}}dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\varphi^{2}\,,$ (31)
@@
3 Thermodynamics
3.1 Thermodynamical quantities
@@
$d\bar{s}_{\rm AdS}^{2}=-\left[1-\frac{1}{3}(1-\ell)\Lambda_{e}r^{2}\right]dt^{2}+\frac{1-\ell}{1-\frac{1}{3}(1-\ell)\Lambda_{e}r^{2}}dr^{2}+r^{2}(d\theta^{2}+r^{2}\sin^{2}\theta d\phi^{2}),$ (37)
@@
[…]
@@ 3.2.1 Case A
Obviously, the value range of $\mathcal{R}$ is uncertain. If $\ell>0$ (namely, $\xi_{2}>0$), then $\mathcal{R}<1$. In this case, the static neutral AdS black hole in KR gravity violates the RII and is superentropic. Otherwise, if $\ell\leq 0$, one then has $\mathcal{R}\leq 1$. In this situation, the static neutral AdS black hole in KR gravity obeys the RII and is subentropic. Therefore, the static neutral AdS black hole in KR gravity can obey or violate the RII, and it is either subentropic or superentropic, depending upon the range of the coupling constant $\xi_{2}$. This character is very similar to that of the ultraspinning (dyonic) Kerr-Sen-AdS4 black holes Wu et al. (2020, 2021) and the six-dimensional ultraspinning Chow’s black holes in gauged supergravity Wu and Wu (2021).
3.2.2 Case B

### v2

# Static neutral black holes in Kalb-Ramond gravity

@@ Abstract
I Introduction
The search for quantum gravity effects has attracted attention over the last decades. On the experimental side, we have witnessed remarkable advancements, such as gravitational wave detections produced by LIGO and Virgo Abbott et al. (2016a, b, 2017, 2020), as well as high-energy particle collisions at the LHC Aad et al. (2012); Chatrchyan et al. (2012); Aaij et al. (2023). On the theoretical side, the study of Lorentz symmetry breaking (LSB) is important for understanding quantum gravity processes in fundamental physics Kostelecky and Potting (1991); Colladay and Kostelecky (1998); Kostelecky and Samuel (1989a); Gambini and Pullin (1999); Carroll et al. (2001); Tian and Du (2021); Tian et al. (2022); PRD109-026004 . By investigating the low-energy contributions from Lorentz symmetry breaking, especially its impact on spacetime, and analyzing high-precision experimental data, we can explore the possibility of Lorentz symmetry breaking in spacetimes compared to general relativity (GR). This analysis also allows us to assess whether these implications are consistent with a quantum gravity model, providing an accurate description of the observed phenomena within that framework.
@@ I Introduction
Another significant Lorentz violation model is the Kalb-Ramond (KR) gravity theory, which incorporates a nonminimally coupled tensor field $B_{ab}$, known as the KR field. This field is a tensorial field arising from the bosonic spectrum of string theory Kalb and Ramond (1974); Kao et al. (1996). It is an antisymmetric 2-tensor, denoted as $B_{[ab]}$, and can be conveniently decomposed as $B_{ab}=\tilde{E}_{[a}v_{b]}+\epsilon_{abcd}v^{c}\tilde{B}^{d}$, where $v^{a}$ is a timelike 4-vector. The pseudofields $\tilde{E}_{a}$ and $\tilde{B}_{a}$ are spacelike, satisfying $\tilde{E}_{a}v^{a}=\tilde{B}_{a}v^{a}=0$. Analogous to Maxwell electrodynamics, these fields can be interpreted as the pseudoelectric and pseudomagnetic fields, respectively. Therefore, the KR VEV yields two background vectors, in contrast to the single vector produced by the bumblebee VEV. In this framework, considerable research has been conducted on finding exact solutions in the Einstein-Kalb-Ramond field equations Chakraborty and SenGupta (2016); Maluf and Muniz (2022); Yang et al. (2023); Duan et al. (2023). Notably, the pioneering work by Yang et al. Yang et al. (2023) first correctly derived the Schwarzschild-like solution in this theory, and the Schwarzschild-(A)dS-like solution was also obtained by relaxing the vacuum conditions. Subsequently, the study of spacetime properties surrounding these black hole solutions has garnered significant attention in the past year, including research on black hole quasi-normal mode (QNM) frequencies Guo et al. (2023); Filho et al. (2024a), photon spheres and shadows Jha (2024); Junior et al. (2024a), gravitational lensing Junior et al. (2024b); 2406.11582 , geodesic motion Filho et al. (2024b), etc Du et al. (2024); PRD105-103505 . However, these solutions were limited to the specific spherically symmetric scenario $(-g_{tt}=g_{rr}^{-1})$ and did not extend to the more general spherically symmetric black hole scenario $(-g_{tt}\neq g_{rr}^{-1})$. To achieve this goal, it is essential to obtain a comprehensive understanding of solutions for static neutral spherically symmetric black holes solutions in KR gravity theory.
@@
The remaining part of this paper is organized as follows. In Sec. II, we provide a concise overview of the Kalb-Ramond gravity theory. We solve the equations of the theory to obtain black hole exact solutions for two types of static neutral spacetimes (the cases where $-g_{tt}=g_{rr}^{-1}$ and $-g_{tt}\neq g_{rr}^{-1}$), considering both with and without the cosmological constant. In Sec. III, some basic thermodynamic properties of the obtained Kalb-Ramond black holes are analyzed. In Sec. IV, we explore the dynamical evolution of the scalar field for both types of spacetimes. Both types of solutions decay rapidly with an increase in the Lorentz-violating parameter, with the latter type being more sensitive. Finally, our conclusions and outlooks are given in Sec. V.
II KR black hole solutions
@@
$T^{\text{KR}}_{ab}=$ $\frac{1}{2}H_{acd}H_{b}{}^{cd}-\frac{1}{12}g_{ab}H^{cde}H_{cde}+2V^{\prime}(X)B_{ca}B^{c}{}_{b}$ (3)
$-g_{ab}V(X)+\xi_{2}\bigg{[}\frac{1}{2}g_{ab}B^{ce}B^{d}{}_{e}R_{cd}-B^{c}{}_{a}B^{d}{}_{b}R_{cd}$
$-B^{cd}B_{bd}R_{ac}-B^{cd}B_{ad}R_{bc}+\frac{1}{2}\nabla_{c}\nabla_{a}(B^{cd}B_{bd})$
$+\frac{1}{2}\nabla_{c}\nabla_{b}(B^{cd}B_{ad})-\frac{1}{2}\nabla^{c}\nabla_{c}(B_{a}{}^{d}B_{bd})$
$-\frac{1}{2}g_{ab}\nabla_{c}\nabla_{d}(B^{ce}B^{d}{}_{e})\bigg{]},$
which can be considered as the energy-momentum tensor of the KR field. The prime indicates the derivative with respect to the variable of the corresponding functions. On the other hand, by varying the action (1) with respect to the KR field $B_{ab}$, we can obtain the equation of motion for the KR field, which is given by $\Pi_{ab}=0$, where
$\Pi_{ab}=\nabla^{c}H_{abc}+3\xi_{2}R_{ca}B^{c}_{~{}b}-3\xi_{2}R_{cb}B^{c}_{~{}a}-6V^{\prime}B_{ab}.$ (4)
@@
it follows that the vacuum field exhibits a pseudoelectric configuration. Therefore, this configuration automatically causes the KR field strength to vanish, i.e., $H_{abc}=0$ or $H_{3}=d\beta_{2}=0$ Yang et al. (2023). Regarding the potential $V$, a prime example that meets these conditions is a smooth quadratic function, similar to the one proposed by Casana et al. Casana et al. (2018), given by:
$V=V(X)=\frac{1}{2}\lambda X^{2},$ (5)
where $\lambda$ is a constant and $X$ represents a generic potential argument. Consequently, the VEV, $\beta_{ab}$, arises as a solution of $V=V^{\prime}=0$. However, this assumption is only applicable in the special case where the cosmological constant is not considered. To obtain asymptotically (A)dS spacetime in this theory, that is, when a nonzero cosmological constant exists, consider relaxing the vacuum conditions. Similar to the potential given by Maluf et al. Maluf and Neves (2021), another simple choice is a linear function defined as:
$V=V(\lambda,X)=\lambda X.$ (6)
Here, $\lambda$ is interpreted as a Lagrange-multiplier field Bluhm et al. (2008). Its equation of motion ensures the vacuum condition $X=0$, leading to $V=0$ for any field $\lambda$ on-shell. Notably, this form of potential manifests itself as a cosmological constant. Thus, the potential and its first-order derivative satisfy
$V(\beta^{ab}\beta_{ab}+b^{2})=\lambda(\beta^{ab}\beta_{ab}+b^{2})=0,$ (7)
$V^{\prime}(\beta^{ab}\beta_{ab}+b^{2})=\lambda,$ (8)
where $V^{\prime}(X)=dV(X)/dX$. Then, we can define the efficient gravitational field equation, satisfying $\mathcal{G}_{ab}=0$, as follows:
$\mathcal{G}_{ab}=R_{ab}-\Lambda g_{ab}-\xi_{2}\mathcal{B}_{ab}-\lambda(2\beta_{ac}\beta_{b}{}^{c}+b^{2}g_{ab}),$ (9)
@@
$\mathcal{B}_{ab}=$ $g_{ab}\beta^{ce}\beta^{d}{}_{e}R_{cd}-\beta^{c}{}_{a}\beta^{d}{}_{b}R_{cd}-\beta^{cd}\beta_{ad}R_{bc}$ (10)
$-\beta^{cd}\beta_{bd}R_{ac}+\frac{1}{2}\nabla_{c}\nabla_{a}(\beta^{cd}\beta_{bd})$
@@
and $\mathcal{G}_{\varphi\varphi}=\mathcal{G}_{\theta\theta}\sin^{2}\theta$, where $\varUpsilon\equiv(A^{\prime}/A)^{2}-2A^{\prime\prime}/A+A^{\prime}B^{\prime}/(AB)$ and $\ell\equiv\xi_{2}b^{2}/2$. The nonzero components of the equations of motion for the KR field are give by
$\Pi_{tr}=6\lambda\tilde{E}(r)+\frac{3\xi_{2}\tilde{E}(r)}{rB}\left[A^{\prime}/A-B^{\prime}/B-r\varUpsilon/2\right],$ (16)
and $\Pi_{rt}=-\Pi_{tr}$. Given $\Pi_{ab}=0$, we can thus find that the specific form of the pseudoelectric field $\tilde{E}(r)$ does not affect the equations of motion for the KR field. It is worth noting that if we consider the Casana condition (5) with $\Lambda=0$, the non-zero components of the KR field equations can be obtained from the combinations of the effective gravitational field equations, as shown below:
$\Pi_{tr}=\frac{3\xi_{2}\tilde{E}(r)}{1-\ell}\left(\mathcal{G}_{tt}/A-\mathcal{G}_{rr}/B\right).$ (17)
Even considering the Maluf condition (7) with $\Lambda\neq 0$, the equation (17) still holds when $\Lambda=(1-\ell)\lambda/\xi_{2}$, which is a necessary condition for this theory to handle asymptotically (A)dS cases Yang et al. (2023). Therefore, in the following discussion, we can focus solely on the gravitational field equations.
II.1 Asymptotically flat spacetime
We first consider asymptotically flat spacetimes, which do not include a cosmological constant, and aim to construct two different types of static neutral spherically symmetric black hole solutions within KR gravity. In this case, we need to assume both the cosmological constant $\Lambda=0$ and the Lagrange-multiplier field $\lambda=0$. Then, using Eqs. (13) and (14), we can obtain a simple relation:
@@ II.1 Asymptotically flat spacetime
$ds^{2}_{\text{A}}=$ $-\left(1-\frac{2M}{r}\right)dt^{2}+\frac{1-\ell}{\left(1-\frac{2M}{r}\right)}dr^{2}$ (23)
$+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\varphi^{2}.$
It is straightforward to verify that the above solution (23) satisfies all field equations (9) when $\Lambda=\lambda=0$. The dimensionless parameter $\ell$ represents the impact of Lorentz violation on spacetime, which is caused by the VEV of the KR field. Note that because the Riemann tensor is given by
@@
$ds^{2}_{\text{B}}=$ $-\left(\frac{1}{1-\ell}-\frac{2M}{r}\right)dt^{2}+\frac{1}{\frac{1}{1-\ell}-\frac{2M}{r}}dr^{2}$ (27)
$+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\varphi^{2}.$
@@
II.2 Asymptotically (A)dS spacetime
If we are to consider the effects of the cosmological constant, we must also consider a non-vanishing Lagrange-multiplier field. In this cases, it is evident that the relation given by Eq. (19) still holds. By setting the metric function $B(r)=\mathcal{C}_{1}A(r)^{-1}$, the equation induced by the relation $\mathcal{G}_{\theta\theta}=0$ is rewritten as
@@ II.2 Asymptotically (A)dS spacetime
$ds^{2}=$ $-\left[1-\frac{2M}{r}-\frac{(1-\ell)\Lambda_{e}}{3}r^{2}\right]dt^{2}$ (32)
$+\frac{(1-\ell)dr^{2}}{1-\frac{2M}{r}-\frac{(1-\ell)\Lambda_{e}}{3}r^{2}}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\varphi^{2}.$
@@
$ds^{2}=$ $-\left[\frac{1}{1-\ell}-\frac{2M}{r}-\frac{\Lambda_{e}}{3}r^{2}\right]dt^{2}$ (34)
$+\frac{1}{\frac{1}{1-\ell}-\frac{2M}{r}-\frac{\Lambda_{e}}{3}r^{2}}dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\varphi^{2}\,,$
@@
III Thermodynamics
III.1 Thermodynamical quantities
@@
$d\bar{s}_{\rm AdS}^{2}=$ $-\left[1-\frac{1}{3}(1-\ell)\Lambda_{e}r^{2}\right]dt^{2}+\frac{1-\ell}{1-\frac{1}{3}(1-\ell)\Lambda_{e}r^{2}}dr^{2}$
$+r^{2}(d\theta^{2}+r^{2}\sin^{2}\theta d\phi^{2}),$ (40)
@@
$\mathcal{Q}[\chi]=$ $\frac{1}{16\pi}\int_{0}^{2\pi}d\varphi\int_{0}^{\pi}d\theta\sqrt{-\bar{g}}\big{(}\chi_{c}{\nabla}_{d}\mathcal{K}^{trcd}$ (43)
$+\mathcal{K}^{tcdr}{\nabla}_{d}\chi_{c}\big{)}|_{r\to\infty},$
@@ III.2.1 Case A
Obviously, the value range of $\mathcal{R}$ is uncertain. If $\ell>0$ (namely, $\xi_{2}>0$), then $\mathcal{R}<1$. In this case, the static neutral AdS black hole in KR gravity violates the RII and is superentropic. Otherwise, if $\ell\leq 0$, one then has $\mathcal{R}\geq 1$. In this situation, the static neutral AdS black hole in KR gravity obeys the RII and is subentropic. Therefore, the static neutral AdS black hole in KR gravity can obey or violate the RII, and it is either subentropic or superentropic, depending upon the range of the coupling constant $\xi_{2}$. This character is very similar to that of the ultraspinning (dyonic) Kerr-Sen-AdS4 black holes Wu et al. (2020, 2021) and the six-dimensional ultraspinning Chow’s black holes in gauged supergravity Wu and Wu (2021).
III.2.2 Case B